OpenStudy (anonymous):

A sequence is defined by the formula f(n + 1) = f(n) – 3. If f(4) = 22, what is f(1)?

OpenStudy (welshfella):

f(3 +1) = f(3) - 3 f(4) - f(3) = -3 22 - f(3) = -3 f(3) = 22+3 = 25 so f(2) = 25 + 3 = 28 and f(1) = ?

OpenStudy (anonymous):

28?

OpenStudy (welshfella):

No f(2) = 28

OpenStudy (welshfella):

do you see the pattern the terms are going down 3 as we start form 1 ? 28,25,22 etc so what would be the first term?